# TV-TREES: Multimodal Entailment Trees for Neuro-Symbolic Video Reasoning

### v3

@@ Abstract
It is challenging to perform question-answering over complex, multimodal content such as television clips. This is in part because current video-language models rely on single-modality reasoning, have lowered performance on long inputs, and lack interpretability. We propose TV-TREES, the first multimodal entailment tree generator. TV-TREES serves as an approach to video understanding that promotes interpretable joint-modality reasoning by producing trees of entailment relationships between simple premises directly entailed by the videos and higher-level conclusions. We then introduce the task of multimodal entailment tree generation to evaluate the reasoning quality of such methods. Our method’s experimental results on the challenging TVQA dataset demonstrate interpretable, state-of-the-art zero-shot performance on full video clips, illustrating a best-of-both-worlds contrast to black-box methods.
@@ 1 Introduction
Figure 1: A (a) QA pair and (b) corresponding video clip and dialogue from the TVQA dataset Lei et al. (2018) and (c) a multimodal entailment tree, recursively produced by our approach (top-down). Trees are created by recursively retrieving atomic evidence from the transcript and video frames and decomposing the QA pair into compositionally equivalent hypotheses until each can be directly entailed by the retrieved evidence.
@@
Videos account for a large portion of content available and consumed online, but automated reasoning over semantically complex video-language data remains a challenging and under-explored problem. A popular task for assessing models’ video understanding is narrative-centric video question-answering (VideoQA): Given a natural language question, a video clip of a movie or TV show, and a corresponding dialogue transcript, the goal is to return a correct natural language answer to the question using the video-text data.
@@
Methods tackling this task Yang et al. (2022); Li et al. (2020); Ko et al. (2023) frequently take the form of large, joint-modality transformer models. While these systems typically outperform smaller, domain-specific architectures, they inherently lack qualities necessary for robust and reliable video-language understanding. In addition to model performance often correlating with the length of the input video clip, analyses suggest their ability to perform joint visual-language reasoning is also limited, that they rely on either text or visual content but not both Rawal et al. (2023). Better interpretability of these models could illuminate these reasoning pitfalls and allow researchers to identify and correct system issues. However, while LLMs now facilitate increasingly transparent explanation generation alongside outputs Zhao et al. (2023), video-language models lack this ability.
@@
Entailment trees Dalvi et al. (2021), or trees of entailment relationships between atomic premises and higher-level conclusions, have been shown to serve well as the structural basis for text-only QA tasks by systematically and transparently modeling logical reasoning chains Weir and Van Durme (2023). We embrace this approach: We develop (1) the first multimodal entailment tree generator, TV-TREES (the Transparent Video-Text REasoning with Entailment System), and (2) the task of multimodal entailment tree generation to assess the reasoning ability of such systems.
@@
In contrast to existing black-box systems, TV-TREES focuses on the manipulation of atomic "facts" retrieved from video clips to answer VideoQA questions. The approach jointly reasons over both modalities and is compatible with long video inputs, and crucially, the resulting entailment trees provide human-interpretable evidence and natural language explanations for each logical operation. Our evaluation method builds on work in informal logic and textual entailment tree generation, adapting these ideas to the multimodal domain with an emphasis on reliable evaluation.
@@
We show that our multimodal reasoning system performs competitively on zero-shot VideoQA for the difficult TVQA dataset Lei et al. (2018), while at the same time providing interpretable reasoning traces. Further, TV-TREES achieves state-of-the-art performance using full-length video clips as input.
@@
    The first multimodal entailment tree generator, a fully explainable video understanding system that emphasizes logical reasoning across modalities.
@@
    The task of multimodal entailment tree generation and a corresponding metric for evaluating step-by-step video-text reasoning quality.
@@
    Results demonstrating state-of-the-art performance on zero-shot TVQA when using full clips and transcripts as input.
@@ 2.1 VideoQA
QA over images makes up a large portion of multimodal question-answering work Zou and Xie (2020). VideoQA benchmarks constitute a smaller portion of this area Zhong et al. (2022) and often focus on simple content and questions Jang et al. (2017), but some recent VideoQA datasets have targeted models’ commonsense knowledge and inference ability Lei et al. (2018); Zadeh et al. (2019). Recently, vision-and-language transformers have substantially improved performance on these VideoQA tasks Zhou et al. (2020), and can often reason over complex content without an external knowledge base Kim et al. (2021); Wang et al. (2021b); Salin et al. (2022).
In contrast to these video-language models, Khurana and Deshpande (2021) highlight alternative deep learning strategies for VideoQA such as attention-free methods, attention-based methods, memory network methods, and hierarchical reinforced methods. Notably, Zhao et al. (2018, 2020) propose a hierarchical encoder-decoder model that uses adaptive video segmentation based on the question contents. Related works consider graph networks for video understanding Wang et al. (2021a); Gu et al. (2021); Liu et al. (2022). While these models scale to longer videos more successfully, their performance suffers compared to transformer-based approaches.
@@ 2.2 Explainable Multimodal Understanding
Traditional techniques like kernel visualization and perturbation have been considered for video explainability Hiley et al. (2019); Li et al. (2021b) alongside other approaches that consider low-level reasoning steps for simple tasks Zhuo et al. (2019); Roy et al. (2019); Nourani et al. (2020). Some work focuses on grounded VideoQA, in which models are tasked with providing the visual evidence necessary for answering a question about spatial-temporal content Xiao et al. (2023).
The approaches most similar to our work are Chen and Kong (2021) and Mao et al. (2022). Chen and Kong (2021) tackle the VIOLIN video entailment dataset Liu et al. (2020) by grounding the relevant textual entities in the video and transcript and providing a heatmap over the input as an explanation for the produced output. Our work differs in that we show exactly what data pieces contribute to the final output, explicitly model each step of the reasoning process, and don’t require fine-tuning on the target dataset or domain. Mao et al. (2022) uses a chain-of-thought explanation system based on a video scene graph to answer questions about actions and objects in short video clips and GIFs. The primary difference between this and our work is the lack of dialogue and visual semantic complexity. The chain-of-thought reasoning primarily considers logical and taxonomy-centric operations over atomic-level scene graph content instead of complex inference reasoning, and the input for their proposed system only spans a few seconds.
@@ 2.3 Entailment Tree Generation
This paper draws inspiration from recent work on constructing natural language entailment trees to explain reasoning. The notion starts with Dalvi et al. (2021), who introduce an expert-annotated dataset of compositional trees showing how a hypothesis follows as a logical consequence of a series of multi-premise entailment steps starting from verified support facts. They propose a series of reconstruction tasks, challenging models to reproduce expert-annotated trees given just the top-level hypothesis and some amount of gold and distractor fact leaves, and our proposed multimodal construction task is inspired by this formulation.
More recent work has introduced methods to tackle Dalvi et al.’s reconstruction task Bostrom et al. (2022); Neves Ribeiro et al. (2022), and to use entailment trees as a basis for neuro-symbolic reasoning Tafjord et al. (2022); Weir and Van Durme (2023). Our work is most similar to Weir and Van Durme (2023), who introduce a QA system that reasons by searching via backward chaining for entailment trees grounded in a knowledge source. We build upon this notion, extending it to the multimodal setting and addressing the many resulting challenges.
@@ 2.4 Multimodal Entailment
There is a selection of work that considers entailment in images and video: Xie et al. (2019) introduce a dataset of image-entailment pairs similar to the SNLI Bowman et al. (2015a) corpus, and Do et al. (2020) add natural language explanations to the pairs. More specific visual entailment tasks in this domain have been proposed as well Thomas et al. (2022); Li et al. (2023b)., and Suzuki et al. (2019) introduce a logic system for identifying entailment between images and captions.
Notably, Liu et al. (2020) introduce VIOLIN, a dataset of videos paired with natural language inferences that are either entailed or contradicted by the video content. Typically, standard vision-language transformers are trained for this task Li et al. (2020); Sun et al. (2022), but more tailored approaches exist as well Li et al. (2021a); Chen and Kong (2021).
@@ 3 Multimodal Entailment Trees
Figure 2: The multimodal proof tree generator pipeline, matching the contents of Algorithm 1. The dashed boxes divide the pipeline into the three primary modules of the system: The yellow box marks the "retrieval" module, the light blue box marks the "filter" module, and the orange box marks the "decomposition" module. With respect to individual pipeline cells, the light blue and yellow cells represent important pieces of data used or produced during the pipeline, the dark blue cells represent generative text operations, the green cells represent discriminative text operations, and the purple cells represent visual operations.
We introduce the task of multimodal entailment tree generation for the VideoQA domain and the evaluation procedure.
@@ 3.1 Task formulation
Input    Following Dalvi et al. (2021), as input we consider a collection of possible “evidence" and declarative form of a question-answer pair, the hypothesis $h_{(q,a)}$. Traditionally this evidence bank takes the form of a collection of natural language sentences, but in the multimodal domain, it will take the form of a video clip $V$ and corresponding dialogue transcript $D$. The video is an ordered list of $k$ images $V\coloneqq\{v_{i}\}_{i=0}^{k}$, and the transcript is an ordered list of $l$ (dialogue line, timestamp) pairs $D\coloneqq\{(d_{i},s_{i})\}_{i=0}^{l}$, where the timestamp maps the dialogue line to start and end frames within $V$.
Output    We define entailment trees as structures which take the form $T:=(h,e)$. $h$ is a hypothesis and $e$ is evidence, which takes the form of either a
@@
    Leaf: A (possibly empty) subset of items from $V$ or $D$.
@@
    Branch: A pair of two distinct entailment subtrees $T_{1}:=(h_{1},e_{1})$ and $T_{2}:=(h_{2},e_{2})$, where $e:=(T_{1},T_{2})$.
Leaves with empty evidence sets are labeled as null leaves.
The purpose of an entailment tree is to illustrate the compositional reasoning necessary to reach a conclusion from an initial evidence bank through entailment relationships between the parent and child nodes. Therefore, in a well-formed tree, the evidence at any node $(h,e)$ must explicitly entail the hypothesis at that same node. For a leaf node, we posit that $e$ entails $h$ if a human would reasonably infer that $h$ is true if presented only with evidence $e\subseteq V\cup D$. For a branching node, $e$ entails $h$ if a human would reasonably infer that $h$ is true if presented with hypotheses $h_{1}$ and $h_{2}$.
Objective    Given input $(h_{(q,a)},V,D)$, our objective is to return a well-formed entailment tree $T$ that includes null leaves if and only if $a$ is not a correct answer to question $q$.
@@ 3.2 Evaluation
To serve as a secondary and distinct objective from raw VideoQA performance, we propose an evaluation method for assessing the reasoning quality of multimodal entailment trees inspired by Weir et al.’s work on scoring compositional entailments Weir et al. (2024). Informal logic theory posits that natural language arguments may be evaluated in terms of their acceptability, relevance, and sufficiency Johnson and Blair (1977). We consider each node within an entailment tree as an “argument" and consider these qualia as guidelines for comprehensive entailment tree evaluation. Below, we formulate these three qualia through an information-theoretic lens to establish a set of evaluation metrics. We use the Shannon definition of information gain,
@@
where $P(x)$ is the probability that natural language statement $x$ is true conditioned on natural language statement(s) $y$.
Acceptability    Hypotheses at every node should be complete and verifiable natural language statements that are understandable to a human, and hypotheses at leaf nodes should be factually accurate statements conditioned on the world state $(V,D)$. These items may be formalized as
@@
We explore practical implementations of these metrics in Section 5.
@@ 4 TV-TREES
In this section we introduce our proposed multimodal entailment tree generator, beginning with an overview of the framework and then individual module details. All LLM and VLM prompts are included in full in Appendix A.
@@
Figure 3: An example question from TVQA, corresponding dialogue excerpt sampled by TV-TREES, and set of inferences generated from these inputs by TV-TREES. The objective of inference generation is to produce a set of true natural language statements that can help prove the hypothesis.
4.1 System overview
  1. 1.
    Retrieval   Given a hypothesis and a collection of potential evidence, the system first samples relevant evidence from this collection that may sufficiently entail the current hypothesis.
  2. 2.
    Filtering   The system tests whether any retrieved evidence fully entails the hypothesis. If such evidence exists and was retrieved, it is returned and the current node becomes a leaf.
  3. 3.
    Decomposition   If the retrieval and filtering steps result in insufficient evidence, the system decomposes the hypothesis into two sub-hypotheses such that proving both independently is equivalent to proving the original hypothesis.
The interaction of these three parts is illustrated in Algorithm 1. Given a hypothesis $h$, transcript sample $D^{\prime}\subseteq D$, and video sample $V^{\prime}\subseteq V$, the system first returns evidence from the transcript relevant to $h$ (line 1) and identifies whether any of it entails $h$ (2). If such evidence was retrieved, $e$ is set to the best sample (4), and the leaf node is returned (17). Otherwise, $h$ is decomposed into sub-hypotheses $h_{0}$ and $h_{1}$ (8) and the algorithm is recursively called on these newly constructed sub-problems (9-10), treating the generated sub-proofs as explanation $e$ (11). If textual evidence cannot be found for the current node nor any of the downstream nodes (14), then the visual evidence in sample $V^{\prime}$ is sampled, filtered, (13) and assigned to $e$ where applicable (15) in the same manner as the text content.
If the maximum depth is reached during recursion, the evidence at that node is set to the empty set and the tree is incomplete.
In the following sections, we explain the implementation of the subroutines called by Algorithm 1.
@@ 4.2 Preprocessing
Hypothesis Generation    The purpose of the hypothesis generation is to provide the downstream modules with a single declarative statement that contains the full semantic meaning of the original QA pair. For simplicity, this generative operation is carried out by prompting GPT-3.5 Brown et al. (2020). We find that less robust in-context learning models like FLAN-T5 Chung et al. (2022) are prone to omitting contextual details present in the question and not handling typos appropriately.
@@
Evidence Localization    Given the hypothesis, TV-TREES attempts to identify a temporal window to sample evidence based on the dialogue. We use a cross-encoder model trained on the MS MARCO passage ranking task Bajaj et al. (2016) to rank six-line transcript passages on their computed similarity with the generated hypothesis. We use a sliding window to calculate scores for every potential sample and return the highest-scoring excerpt. If a sufficient window is identified, the vision pipeline inherits this window. If no sufficient dialogue sample is found, the system uses all video frames as the evidence bank, omitting text entirely.
4.3 Evidence Retrieval

### v4

@@ Abstract
It is challenging for models to understand complex, multimodal content such as television clips, and this is in part because video-language models often rely on single-modality reasoning and lack interpretability. To combat these issues we propose TV-TREES, the first multimodal entailment tree generator. TV-TREES serves as an approach to video understanding that promotes interpretable joint-modality reasoning by searching for trees of entailment relationships between simple text-video evidence and higher-level conclusions that prove question-answer pairs. We also introduce the task of multimodal entailment tree generation to evaluate reasoning quality. Our method’s performance on the challenging TVQA benchmark demonstrates interpretable, state-of-the-art zero-shot performance on full clips, illustrating that multimodal entailment tree generation can be a best-of-both-worlds alternative to black-box systems.
@@ 1 Introduction
Figure 1: A QA pair and corresponding video clip and dialogue from the TVQA dataset Lei et al. (2018), and a multimodal entailment tree, recursively produced by our approach (top-down). Trees are created by recursively retrieving atomic evidence from the transcript and video frames and decomposing the QA pair into compositionally equivalent hypotheses until each can be directly entailed by the retrieved evidence.
@@
Videos account for a large portion of content available and consumed online, but automated reasoning over semantically complex video-language data remains a challenging and under-explored problem. A popular task for assessing models’ video understanding is narrative-centric video-language question-answering: Given a natural language question, a video clip of a movie or TV show, and a corresponding dialogue transcript, the goal is to return a correct natural language answer to the question using the video-language data.
@@
Methods tackling the video-language QA task Yang et al. (2022); Li et al. (2020); Ko et al. (2023) frequently take the form of large, joint-modality transformer models. Analyses suggest their ability to perform joint visual-language reasoning is limited, as they often rely on either text or visual content but not both Rawal et al. (2023). Assessing modality reliance and reasoning quality overall is difficult given their lack of interpretability: While LLMs now facilitate increasingly transparent explanation generation alongside outputs Zhao et al. (2023), video-language models generally lack this ability.
@@
Entailment trees Dalvi et al. (2021), or trees of entailment relationships between atomic premises and higher-level conclusions, have been shown to serve well as the structural basis for text-only QA tasks by systematically and transparently modeling logical reasoning chains Weir et al. (2024a). We embrace this approach: We develop (1) the first multimodal entailment tree generator, TV-TREES (the Transparent Video-Text REasoning with Entailment System), and (2) the task of multimodal entailment tree generation to assess the reasoning ability of such systems.
@@
In contrast to existing black-box QA systems, TV-TREES focuses on the search and manipulation of atomic “facts” retrieved from video clips to search for proofs for video-language question-answer pairs. The approach jointly reasons over both modalities and, crucially, the resulting entailment trees provide human-interpretable evidence and natural language explanations for each logical operation, enabling direct analysis of the model’s underlying reasoning. Our entailment tree evaluation method builds on work in informal logic, adapting these ideas to the multimodal domain with an emphasis on reliable evaluation.
@@
We show that our multimodal reasoning system performs competitively on zero-shot video-language QA for the difficult TVQA dataset Lei et al. (2018) while simultaneously providing comprehensive and interpretable reasoning traces. Further, TV-TREES achieves state-of-the-art performance using full-length clips as input.
@@
    The first multimodal entailment tree generator, an explainable video-language understanding system that emphasizes logical reasoning across modalities by searching for proofs for question-answer pairs.
@@
    The task of multimodal entailment tree generation for evaluating step-by-step video-language reasoning quality.
@@
    Results demonstrating state-of-the-art zero-shot video-language QA performance on TVQA, ablation experiments demonstrating the benefit of joint-modality reasoning, and quantitative and qualitative analyses of entailment tree quality.
@@ 2.1 VideoQA
QA over images makes up a large portion of multimodal question-answering work Zou and Xie (2020). VideoQA benchmarks constitute a smaller portion of this area Zhong et al. (2022) and often focus on simple content and questions Jang et al. (2017), but some recent VideoQA datasets have targeted models’ commonsense knowledge and inference ability, namely TVQA and MovieQA Lei et al. (2018); Tapaswi et al. (2016). We focus our experiments on TVQA as evidence suggests about half of MovieQA’s questions can be answered with the question and answer options alone. Jasani et al. (2019)
Recently, vision-and-language transformers have substantially improved performance on these VideoQA tasks Zhou et al. (2020), and can often reason over complex content without an external knowledge base Kim et al. (2021); Wang et al. (2021); Salin et al. (2022). In contrast to these video-language models, Khurana and Deshpande (2021) highlight alternative strategies for VideoQA such as attention-free methods, attention-based methods, memory network methods, and hierarchical reinforced methods. Notably, Zhao et al. (2018, 2020) propose a hierarchical encoder-decoder model that uses adaptive video segmentation based on the question contents.
@@ 2.2 Explainable Multimodal Understanding
Traditional techniques like kernel visualization and perturbation have been considered for video explainability Hiley et al. (2019); Li et al. (2021b) alongside other approaches that consider low-level reasoning steps for simple tasks Zhuo et al. (2019); Roy et al. (2019); Nourani et al. (2020). Additionally, Lu et al. (2022) introduce a transparent reasoning benchmark for vision-text QA. The approaches most similar to our work are Chen and Kong (2021) and Mao et al. (2022). Chen and Kong (2021) ground relevant textual entities in video and dialogue through a heatmap over the input as an explanation for the produced output. Our work differs in that we show exactly what data pieces contribute to the final output, explicitly model each step of the reasoning process, and don’t require fine-tuning on the target dataset or domain. Mao et al. (2022) uses a chain-of-thought explanation system based on a video scene graph to answer questions about actions and objects in short video clips and GIFs. This method does not consider dialogue and focuses on simple visual questions, instead of complex inferential reasoning that TV-TREES tackles. Furthermore, the input for their proposed system only spans a few seconds.
@@ 2.3 Entailment Tree Generation
This paper draws inspiration from recent work on constructing natural language entailment trees to explain reasoning. The notion starts with Dalvi et al. (2021), who introduce an expert-annotated dataset of compositional trees showing how a hypothesis follows as a logical consequence of a series of multi-premise entailment steps starting from verified support facts. More recent work has introduced methods to tackle Dalvi et al.’s reconstruction task Bostrom et al. (2022); Neves Ribeiro et al. (2022) and to use entailment trees as a basis for neuro-symbolic reasoning Tafjord et al. (2022); Weir et al. (2024a). Our work is most similar to Weir et al. (2024a), who introduce a QA system that reasons by searching for entailment trees grounded in a knowledge source. We extend this notion to the multimodal setting and address the resulting challenges.
@@ 2.4 Multimodal Entailment
There is a selection of work that considers entailment in images and video: Xie et al. (2019) introduce a dataset of image-entailment pairs similar to the SNLI Bowman et al. (2015a) corpus, and Do et al. (2020) add natural language explanations to the pairs. More specific visual entailment tasks in this domain have been proposed as well Thomas et al. (2022); Li et al. (2023b)., and Suzuki et al. (2019) introduce a logic system for identifying entailment between images and captions. Liu et al. (2020) introduce VIOLIN, a dataset of videos paired with natural language inferences that are entailed or contradicted by the video content, and traditional models Li et al. (2020); Sun et al. (2022) as well as tailored approaches Li et al. (2021a); Chen and Kong (2021) are trained for the task.
@@ 3 Multimodal Entailment Trees
Figure 2: The TV-TREES pipeline. The system searches for evidence in the video clip and transcript that it can use to prove higher-level conclusions, with the goal of recursively constructing a tree of entailment relationships between these conclusions and low-level evidence. The figure highlights the system’s three primary modules: Retrieval, filtering, and decomposition. Retrieval involves identifying the best time interval of the video to sample from and then extracting natural language inferences and video frames from the input data. Filtering involves filtering the extracted data samples with NLI classifiers and VQA systems to identify evidence that proves the answer to the question. Decomposition addresses when no evidence is found— an LLM is used to decompose the question and hypothesis into two smaller sub-claims, to be each recursively proven through the same process.
We now introduce the task of multimodal entailment tree generation for video-language QA.
@@ 3.1 Task formulation
Input    Following Dalvi et al. (2021), as input we consider hypothesis $h_{(q,a)}$ (the declarative form of a question-answer pair) and an evidence bank. Traditionally, the evidence bank takes the form of a set of natural language sentences, but in the multimodal domain it is a video clip $V$ and corresponding dialogue transcript $D$, treated as sets of video frames and dialogue lines, respectively.
Output
We define entailment trees as recursive structures which take the form $T:=(h,e)$. $h$ is a hypothesis and $e$ is evidence. $e$’s form is either a
@@
    Leaf: A (possibly empty) subset of items from evidence bank $\{V\cup D\}$.
@@
    Branch: A pair of two distinct entailment subtrees $T_{1}:=(h_{1},e_{1})$ and $T_{2}:=(h_{2},e_{2})$.
Leaves with empty evidence sets are labeled null.
The purpose of an entailment tree is to illustrate the compositional reasoning necessary to reach a conclusion from an initial evidence bank using entailment relationships. In a well-formed tree, the evidence $e$ in any tree node $(h,e)$ must explicitly entail the hypothesis $h$. For a leaf node, $e$ entails $h$ if a human would reasonably infer that $h$ is true given evidence $e\subseteq\{V\cup D\}$. For a branching node, $e$ entails $h$ if a human would reasonably infer that $h$ is true given hypotheses $h_{1}$ and $h_{2}$.
Objective
Given inputs $h_{(q,a)},V,\text{ and }D$, our objective is to return a well-formed entailment tree $T$ that includes null leaves if and only if $a$ is not a correct answer to question $q$.
@@ 3.2 Evaluation
Acceptability
Hypotheses at every node should be complete and verifiable natural language statements that are understandable to a human, and hypotheses at leaf nodes should be factually accurate statements conditioned on the world state $\{V\cup D\}$. These items may be formalized as
@@
Given these metric formulations, we explore practical implementations of them in Section 5.
@@ 4 TV-TREES
We now introduce our multimodal entailment tree generator, pictured in Figure 2.
4.1 System overview
  1. 1.
    Retrieval   Given a hypothesis and evidence bank, the system samples evidence candidates from the bank that may sufficiently entail the current hypothesis.
  2. 2.
    Filtering   The system tests whether any retrieved evidence entails the hypothesis. If such evidence was retrieved, it is returned and the current tree node becomes a leaf.
  3. 3.
    Decomposition   If the previous steps result in insufficient evidence, the system decomposes the hypothesis into two sub-hypotheses such that proving both independently is equivalent to proving the original hypothesis. The process is recursively called using these sub-hypotheses.
The interaction of these three parts is illustrated in Algorithm 1. Given a hypothesis $h$, transcript sample $D^{\prime}\subseteq D$, and video sample $V^{\prime}\subseteq V$, the system first returns evidence from the transcript relevant to $h$ (line 1) and identifies whether any of it entails $h$ (2). If such evidence was retrieved, $e$ is set to the best evidence (3-4), and the leaf node is returned (17). Otherwise, $h$ is decomposed into sub-hypotheses $h_{0}$ and $h_{1}$ (8) and the algorithm is recursively called on these newly constructed sub-problems (9-10), treating the generated sub-proofs as explanation $e$ (11). If the maximum depth is reached during recursion, the evidence at that node is set to the empty set (5-6). If textual evidence cannot be found for the current node nor any of the downstream nodes (14), then the visual evidence in sample $V^{\prime}$ is sampled, filtered, (13) and assigned to $e$ where applicable (15) in the same manner as the text content. Below, we detail the implementation of the subroutines called by Algorithm 1.
@@ 4.2 Preprocessing
Hypothesis Generation    We first prompt GPT-3.5 Ouyang et al. (2022) to generate a single declarative statement that contains the full semantic meaning of an initial QA pair. [^1] All LLM and VLM prompts are included in Appendix A.
Evidence Localization
@@
Given the hypothesis, TV-TREES attempts to identify a relevant window of the video clip and transcript to sample evidence from. We use a cross-encoder model trained on the MS MARCO passage ranking task Bajaj et al. (2016)222huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 to rank six-line transcript passages on their computed similarity with the generated hypothesis. We use a sliding window to calculate scores for every potential sample and return the highest-scoring excerpt. If a sufficient window is identified, the vision pipeline inherits this same window. If no dialogue sample is found, the system uses all video frames as evidence and omits text entirely.
Figure 3: An example question from TVQA, corresponding dialogue excerpt sampled by TV-TREES, and set of inferences generated from these inputs.
4.3 Evidence Retrieval (Retrieve)